Can someone please help me.
Sum of 2 numbers is 71. The second is 6 more than 4 times the first. What are the 2 numbers.
Thanks

Answer by jayanthihemkos@yahoo.com(87)

Sum of 2 numbers is 71. The second is 6 more than 4 times the first. What are the 2 numbers.

Let first number = x
The second number = 4x+6
THe sum of the two numbers = 71
x+(4x+6) = 71
x+4x+6 = 71
5x+6 = 71
5x+6-6 = 71-6
5x = 65
5x/5 = 65/5
x = 13
CHECK
x+4x+6 = 71
13+4(13) +6 =71
71 = 71
LHS = RHS
Hence x = 13 is the solution